# Scaling Laws for Precision

### v1

@@ Appendix C Alternative Functional Forms
We make these fit choices on sweeps of the form $L(N,D,P_{\text{W}})$ and discuss alternatives to the decomposition/factorization to account for activations and KV cache in Appendix Section K, which assumes an effective parameter count formulation. In this section, a power law refers to a term of the form $C_{\text{w}}\cdot P^{-\alpha_{\text{w}}}$ where $C_{\text{w}},\alpha_{\text{w}}$ are fitted. In general, we find modeling precision effects with power law fits on their own causes the fitted constants $A,B$ to blow up, whereas this does not happen with exponential fits, suggesting the power law does not change sharply enough to match the change in loss induced by precision. We note that while fitting parameters using a double notion of effective parameters and effective data leads to a slightly better fit, it requires more fitted parameters so we stick with the $N_{\text{eff}}$ formulation for simplicity and interpretability. When choosing between fits we validate on held-out data and the $R^{2}$ values below reflect the fit on the held out data. This is in contrast to our plots in the main text, where we have chosen a functional form and we fit and plot on the same data, as is standard in scaling laws [Muennighoff et al., 2024b].
@@ Appendix D Quantization Implementation Details and Types
D.1 Integer Quantization
@@
for channel-based scaling. After scaling, the result is rounded to the nearest integer and then clamped to the range $[Q_{n},Q_{p}]$. After matrix multiplication, the result is rescaled back into the original range. We use integer quantization throughout to fit our scaling laws for simplicity. However, when making pretraining predictions, we test our scaling laws using floating-point quantization since that is used for pretraining in practice.
@@ D.2 Floating-Point Quantization
where $s$ is the sign bit, $m$ is the mantissa, and $e$ is the exponent. In floating-point quantization, both the mantissa and exponent are quantized to reduce the bit width. For exponent-mantissa allocations of bits and details of exponent bias, we follow the guidelines from [Micikevicius et al., 2022] in terms of details, since that reflects how, for instance, FP8 training is done as is common in production settings. Further in line with [Micikevicius et al., 2022], we quantize weights per channel and activations per-tensor.
Making a full scaling law for floating-point quantization is more involved than our integer treatment, because the effects of scaling mantissa vs exponent bits are not the same. In contrast, in integer quantization, each additional bit simply causes us to round into a finer-grained lattice after rescaling, thereby reducing quantization error by a predictable amount. In floating-point quantization, altering the exponent affects the dynamic range, while altering the mantissa changes the precision within that range. This flexibility at once makes floating-point quantization more suitable for model training, but harder to analyze. We leave a commensurately detailed analysis of mantissa vs exponent scaling to future work.
@@ D.3 Hardware Details
Weight-only quantization can accelerate inference because software can be written to accommodate moving data between GPU parts (HBM-SRAM) in smaller units (types), so that a given bandwidth can move more data per second. This reduces memory (IO) bottlenecks that often dominate during inference, even with high-batch workloads. However, we emphasize that the type and therefore speed at which the GPU can do matrix multiplications in natively is determined by the hardware provider, so that even when $P_{\text{w}}=P_{\text{a}}=P_{\text{qkv}}$ (including queries), compute savings are only achieved when these correspond with both a bit-width and type that the GPU supports. For instance, many GPUs support FP32, BF16, INT8, INT4, and the Hopper line of GPUs from NVidia support FP8 matrix multiplication, and the next generation at the time of writing (Blackwell) will support FP4 matrix multiplication. We aim to study scaling in a fairly hardware-agnostic manner so that our work may be useful in the future. We train all our models with fake (simulated) quantization on NVidia H100 GPUs.
@@ Appendix F Replicating PTQ Scaling with other Quantization Methods
Appendix G Why do language models get more sensitive with overtraining?
@@
Appendix H Main Figure Details
@@
Appendix I Numerical Fits
@@
Appendix J Are Weights, Activations, and KV Cache Equally Sensitive?
We find that training runs with $P_{a}\leq 3$ or $P_{\text{kv}}\leq 3$ are not numerically stable, and often diverge, while $P_{\text{w}}=3$ is still well behaved. In particular, we find activations are more sensitive, though this could be because we quantize activations per-tensor and weights-per channel following [Micikevicius et al., 2022], rather than activations being inherently more sensitive. Consequently, we do not fit or validate on runs with activations or attention bits equal to 3. We leave a more detailed analysis of fine-grained sensitivity across layers and types of parameters to future work. The Figure below illustrates the empirical sensitivity by plotting $L(P)$ for the three quantities for various runs $(N,D)$.
@@
Appendix K Empirical $N_{\text{eff}}$
@@
Then plotting this quantity using our fitted numerical values (See Appendix I) gives us the empirical tradeoff between precision and parameters. We can see that the tradeoff is quickly saturating in $P$ to a value near 1. While the functional form is the same for the three model parts, the fitted constants are different. For instance, runs with $P_{\text{a}}\leq 3$ or $P_{\text{kv}}\leq 3$ often diverged, and this was not the case with weight precision. Further, we can see that the KV cache is not sensitive to quantization at higher bit value, but very quickly becomes sensitive around 4-5 bit precision.
@@
Appendix L Additional Plots

### v2

@@ Appendix C Alternative Functional Forms
We make these fit choices on sweeps of the form $L(N,D,P_{\text{W}})$ and discuss alternatives to the decomposition/factorization to account for activations and KV cache in Appendix Section M, which assumes an effective parameter count formulation. In this section, a power law refers to a term of the form $C_{\text{w}}\cdot P^{-\alpha_{\text{w}}}$ where $C_{\text{w}},\alpha_{\text{w}}$ are fitted. In general, we find modeling precision effects with power law fits on their own causes the fitted constants $A,B$ to blow up, whereas this does not happen with exponential fits, suggesting the power law does not change sharply enough to match the change in loss induced by precision. We note that while fitting parameters using a double notion of effective parameters and effective data leads to a slightly better fit, it requires more fitted parameters so we stick with the $N_{\text{eff}}$ formulation for simplicity and interpretability. When choosing between fits we validate on held-out data and the $R^{2}$ values below reflect the fit on the held out data. This is in contrast to our plots in the main text, where we have chosen a functional form and we fit and plot on the same data, as is standard in scaling laws [Muennighoff et al., 2024b].
@@ Appendix D Quantization Implementation Details and Types
D.1 Integer Quantization and Implementation Details
@@
for channel-based scaling. After scaling, the result is rounded to the nearest integer and then clamped to the range $[Q_{n},Q_{p}]$. After matrix multiplication, the result is rescaled back into the original range. We quantize only the forward pass in this work, to ensure fair comparison between quantization-aware-training (weights only) and low-precision training (weights, activations, KV cache). This is because the backward pass is not usually quantized during quantization-aware-training [Ma et al., 2024], so comparing sensitivities of weights (forward only) to activations/KV cache (forward and backward) would not be a principled comparison. In production pretraining in low precision, the matrix multiplications on the backward pass are also quantized, leading to further compute savings. We leave a detailed analysis of how our observations change when accounting for the backward pass to future work. We use integer quantization throughout to fit our scaling laws for simplicity.
@@ D.2 Floating-Point Quantization
where $s$ is the sign bit, $m$ is the mantissa, and $e$ is the exponent. In floating-point quantization, both the mantissa and exponent are quantized to reduce the bit width. For exponent-mantissa allocations of bits and details of exponent bias, we follow the guidelines from [Micikevicius et al., 2022] and quantize weights per channel and activations per-tensor.
Making a full scaling law for floating-point quantization is more involved than our integer treatment, because the effects of scaling mantissa vs exponent bits are not the same. In contrast, in integer quantization, each additional bit simply causes us to round into a finer-grained lattice after rescaling, thereby reducing quantization error by a predictable amount. In floating-point quantization, altering the exponent affects the dynamic range, while altering the mantissa changes the precision within that range. This flexibility at once makes floating-point quantization more suitable for model training, but harder to analyze. We leave a commensurately detailed analysis of mantissa vs exponent – and more generally floating point – scaling to future work.
@@ D.3 Hardware Details
Weight-only quantization can accelerate inference because software can be written to accommodate moving data between GPU parts (HBM-SRAM) in smaller units (types), so that a given bandwidth can move more data per second. This reduces memory (IO) bottlenecks that often dominate during inference, even with high-batch workloads. However, we emphasize that the type and therefore speed at which the GPU can do matrix multiplications in natively is determined by the hardware provider, so that even when $P_{\text{w}}=P_{\text{a}}=P_{\text{qkv}}$ (including queries), compute savings are only achieved when these correspond with both a bit-width and type that the GPU supports. We aim to study scaling in a fairly hardware-agnostic manner so that our work may be useful in the future, and make no claims about hardware details or optimality. We train all our models with fake (simulated) quantization on NVidia H100 GPUs to remain hardware agnostic, not taking advantage of any true low-precision computation. The only assumption is that when hardware does implement support for integer quantization, it is done in a way that involves some combination of rescaling and rounding, as is standard at the time of writing [Dettmers and Zettlemoyer, 2023, Dettmers et al., 2022, Wu et al., 2020, Jacob et al., 2018].
@@ E.2.3 Minimization over $N$, $D$, $P$ with Fixed Compute
E.3 Inference-time Cost Model
For many, inference is the primary cost of training and serving models. Here, we present a preliminary analysis of an inference-time cost model. The key tension is that inference cost scales as $NP$, so that inference costs at a fixed pretraining loss can be reduce by either reducing model size (and overtraining more) or quantizing post-training
We will assume here that $P=P_{\text{post}}$ refers to the precision weights will be quantized to. In practice, inference costs may depend on the precision of the KV cache and activations to some extent as well, but we assume this for tractability of the following mathematical model, and to get a sense of how overtraining and post-train quantization concerns play out at inference-time. We can phrase this minimization problem in the following way.
$\min_{N,D,P}L(N,D,P)=AN^{-\alpha}+BD^{-\beta}+C_{T}\frac{D^{\gamma_{D}}}{N^{\gamma_{N}}}e^{-P/\gamma}\text{ subject to }C=NP$ (29)
The system of first-order conditions that results from this constrained optimization problem is not in general tractable analytically, so we solve the above constrained optimization problem for $P^{*}(C),N^{*}(C),D^{*}(C)$ numerically via a simple grid search. We find that $N^{*},D^{*}$ grow as a power law in $C$ while $P^{*}\propto\log C$. The clumping in points is an artifact of the numerics of the grid search; the fitted lines represent the loglinear (left) and loglog (middle, right) trends overall.
Figure 9: Numerically minimizing a model of inference-time costs with respect to $N,D,P$ after accounting for post-train-quantization degradation and its relation to overtraining.
It might be surprising that $D^{*}$ is not taken to infinity since it does not appear in the cost function. The reason for this is because if it was, post-train degradation (the third term) would become large. It might also be surprising that $D^{*}$ changes with compute at all. The reason for this is because, once again, of the third term: as we allow more inference-time compute we use more $N$, and at a larger $N$ we can now tolerate a larger data budget for a given post-train quantization degradation, so being compute-optimal means taking advantage of this and training that larger parameter count on more data.
The intuition for why $P^{*}\sim\log C$ might be as follows. Consider a situation in which $P^{*}$ is independent of compute: the third term will come to be a bottleneck in loss as compute gets larger because $N,D$ are both being scaled as power laws in compute, and eventually the effect of $e^{-P/\gamma}$ will become non-negligible in comparison to the first two terms in the loss function. To continue decreasing loss at this point, we must make this term smaller at a rate commensurate with the other terms, which go as a power law in compute. Since precision is inside the exponential, this can be done by taking $P\sim\log C$. An important thing to note is that since we are ignoring pretraining costs here, the absolute values of predicted $D^{*}$ are much larger than would be realistically possible in any reasonably training regime, where pretraining costs do matter, even if less than inference costs. But the empirical trends in $N^{*},P^{*}$ showcase how overtraining with post-train quantization in mind can outperform vanilla overtraining without accounting for its effects on post-train quantization.
@@ Appendix F Replicating PTQ Scaling with other Quantization Methods
Appendix G PTQ: Learning Rate Schedule Ablation
Here, we ablate our learning rate and schedule to use warmup with linear decay, as opposed to a cosine schedule, to check it is not an artifact of our choice of learning rate schedule. We do so on our 30M model due to compute constraints, finding the degradation with token/parameter ratio persists, as expected. See Figure 12.
Figure 12: Linear LR Schedule Ablation
Appendix H Why do language models get more sensitive with overtraining?
@@
Appendix I Granularity Ablations
Here, we ablate our choice of quantization granularity (per-tensor vs per-channel) compared to the main text, where we do weights per-channel and activations per-tensor. Per-tensor quantization involves keeping one scalar to rescale all values in a tensor into the quantization codebook range, and per-channel means keeping a scalar per channel dimension; therefore, the latter is strictly more expressive and thus incurs lower quantization loss, than the former, at the cost of slightly more memory usage. Here, we ask: is the increased sensitivity of activations a result of them being inherently more sensitive, or due to the per-tensor design choice.
Figure 13: Quantization granularity ablation: all combination of (training weight precision, training activation precision) $\times$ (per-tensor, per-channel). Dashed lines are per-channel and solid are per-tensor.
These results show that activations are generally more sensitive than weights, since their loss penalty at lower precision goes up faster even when granularity is kept fixed across the two. In fact, quantizing activations per-channel is almost as hard as quantizing weights per-tensor. This is consistent with a broad line of work in quantization finding that activations comprise the central difficulty in quantization [Dettmers and Zettlemoyer, 2023, Ma et al., 2024].
Appendix J Main Figure Details
@@ Appendix I Granularity Ablations
Appendix K Numerical Fits
@@
Appendix L Are Weights, Activations, and KV Cache Equally Sensitive?
We find that training runs with $P_{a}\leq 3$ or $P_{\text{kv}}\leq 3$ are not numerically stable, and often diverge, while $P_{\text{w}}=3$ is still well behaved. In particular, we find activations are more sensitive, though this could be because we quantize activations per-tensor and weights-per channel, rather than activations being inherently more sensitive. Consequently, we do not fit or validate on runs with activations or attention bits equal to 3. We leave a more detailed analysis of fine-grained sensitivity across layers and types of parameters to future work. The Figure below illustrates the empirical sensitivity by plotting $L(P)$ for the three quantities for various runs $(N,D)$.
@@
Appendix M Empirical $N_{\text{eff}}$
@@
Then plotting this quantity using our fitted numerical values (See Appendix K) gives us the empirical tradeoff between precision and parameters. We can see that the tradeoff is quickly saturating in $P$ to a value near 1. While the functional form is the same for the three model parts, the fitted constants are different. For instance, runs with $P_{\text{a}}\leq 3$ or $P_{\text{kv}}\leq 3$ often diverged, and this was not the case with weight precision. Further, we can see that the KV cache is not sensitive to quantization at higher bit value, but very quickly becomes sensitive around 4-5 bit precision.
@@
Appendix N Floating-Point Experiments
The key difference between floating point and integer type is that the former allocates some bits to the exponent representation and some to the mantissa, and these bits play different roles, unlike in integer type where every bit plays the same role in making the quantization lattice uniformly more fine-grained. We hypothesize that if exponent and mantissa bits are scaled jointly (ie. increase together as total bit count does), the overall trend will still be predictable with a functional form like ours. To test this, we fit a parametric form like Equation 3 with the constants $A,B,E,\alpha=\beta$ listed in the table. The overall fit results in values of $\gamma_{w}=2.8111$ and an exponent bias of $b=0.1240$, showing the functional form is still a good fit to the data, even for floating point, under reasonably standard bit allocation schemes between mantissa and exponent. On the middle and right, we again re-fit the same parametric form on now specific values of $(N,D)$ and visualize the quality of the resulting predictions.
Figure 16: Fitting an effective parameter form to floating-point precision for weight training. (Left) involves checking quality of fit on 140 training runs in floating point precision for weights during training.
We use bit allocations of E2M0, E3M0, E4M1, E3M2, E4M2, E5M2, and E5M6 for 3, 4, 5, 6, 7, 8, 12 bits, respectively, with one sign bit throughout. Since exponent and mantissa bits play in general different roles (ie. the effect of a bit on loss and dynamics depends a lot on whether it comes from the mantissa or exponent in floating point), we expect our functional form does well here because mantissa and exponent allocations both increase jointly as precision rises, so overall the trends are predictable in a similar way. We check directly the role of the two by sweeping ExM3 and E3Mx directly, confirming this intuition. This suggests one route for making fine-grained fits for general arbitrary ExMy combinations is to decompose the effects of mantissa and weights, for instance a form like $N_{\text{eff}}(P_{\text{w, m}},P_{\text{w, e}},N)$. Since this is not needed for standard bit allocation choices as we can see in Figure 16, we do not delve into this complexity.
Figure 17: Exponent-mantissa bit allocation sweep. We can see the two types of bits have different scaling behavior, but both fit the saturating form where the first few bits reduce loss a lot, with diminishing returns after that.
Appendix O Additional Plots